OpenStudy (anonymous):

Solve for n and select the correct answer. -2(7n + 10) = 7(4 - 2n)

OpenStudy (across):

Distribute the \(-2\) on the LHS and the \(7\) on the RHS.

OpenStudy (anonymous):

okay and then what?

OpenStudy (across):

What did you obtain?

OpenStudy (anonymous):

-14n+10=28-2n

OpenStudy (anonymous):

-14n-20=28-14n this question is wrong... please check the question again...

OpenStudy (anonymous):

its not wrong

OpenStudy (linknissan):

i will ans this question perfectly

OpenStudy (anonymous):

^^without definition

OpenStudy (anonymous):

(-2 x 7n ) + (-2 x 10) = (7 x 4) + (7 x -2n)

OpenStudy (anonymous):

-14n + -20 = 28 + -14n

OpenStudy (anonymous):

-14n-20=28-14n how can anyone find the value of n from it...?? @linknissan @gaara438125

OpenStudy (anonymous):

wouldn't it be a no solution

OpenStudy (anonymous):

 Yeah , thats is wrong..